OpenStudy (anonymous):

COULD YOU HELP ME PLEASE!!! •Share a problem and solution with a family member. •Determine the number of solutions to a linear system. •Solve a linear system using modeling, guess and check, graphing, and/or algebraic methods. •Check the solution to a linear system. •Report on your findings. I dont have anyone to share a problem with could you share 1 with me and work it out with me please?

OpenStudy (anonymous):

5x = 6+3y is this ok

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Let us look more closely at one example: Example: y = 2x+1 is a linear equation: line on a graph will add attachment of graph The graph of y = 2x+1 is a straight line When x increases, y increases twice as fast, hence 2x When x is 0, y is already 1. Hence +1 is also needed So: y = 2x + 1 Here are some example values: x y = 2x + 1 -1 y = 2 × (-1) + 1 = -1 0 y = 2 × 0 + 1 = 1 1 y = 2 × 1 + 1 = 3 2 y = 2 × 2 + 1 = 5 Check for yourself that those points are part of the line above!

OpenStudy (anonymous):

does this help you out any

OpenStudy (anonymous):

do u understand @Ariel-red

OpenStudy (anonymous):

thanks